OpenStudy (love_to_love_you):

Given the function f(x) = x^2 + 13.5x + 45, use Newton's Method to find x⌄6 when x⌄0 = -7.

OpenStudy (love_to_love_you):

the 6 and 0 are supposed to be below the x

OpenStudy (raffle_snaffle):

https://en.wikipedia.org/wiki/Newton's_method I know...

OpenStudy (raffle_snaffle):

You show her/him and I will watch. XD

OpenStudy (royalranger):

Do you even know how to use the formula?

OpenStudy (royalranger):

@love_to_love_you

OpenStudy (raffle_snaffle):

Sort of maybe kind of. It has been awhile... That is why I wanted you to do it. xD

OpenStudy (raffle_snaffle):

@zepdrix

OpenStudy (love_to_love_you):

Nope.

OpenStudy (raffle_snaffle):

@mathmale

OpenStudy (royalranger):

Ok, first off find the derivative of the function, since we will need it later.

OpenStudy (royalranger):

You there?

OpenStudy (royalranger):

Guess not, good luck homie

OpenStudy (raffle_snaffle):

She isn't here... She is gone...

OpenStudy (eliesaab):

To check your answers here are all \[ x_i \text{ for } i=0,1,2,3,4,5,6} \] {-7, -8., -7.6, -7.50588, -7.50002, -7.5, -7.5}

OpenStudy (eliesaab):

You can check that\[ x^2+13.5 x+45=(x+6) (x+7.5) \] so we got the actual root

OpenStudy (mathmale):

Would be clearer for everyone if you'd please use the correct notation: Given the function f(x) = x^2 + 13.5x + 45, use Newton's Method to find x⌄6 when x⌄0 = -7. Confusing Given the function f(x) = x^2 + 13.5x + 45, use Newton's Method to find \[x _{6}\]

OpenStudy (mathmale):

when the initial, "guessed," starting value is \[x _{0}=-7\]

OpenStudy (mathmale):

do you know the formula for Newton's method? If so would you mind typing it out so that we can verify that you're working with the right formula?